# Nevanlinna Theory of Algebroid Functions on Complete Kähler Manifolds

### v10

@@ Abstract.
- 3 Uniformization of Algebroid Functions
@@ 1.1. Motivation
where are holomorphic functions without common zeros in An algebroid function is viewed as a special multi-valued analytic function, while which was first put forward, G. Darboux deemed that it is an important class of functions. To see that, for example, it is known that a -valued algebroid function defines a -leaf Riemann surface which is a generic -sheeted analytic covering of A further fact states that can lift to a single-valued meromorphic function on via a natural projection such that which is called the uniformization of
@@ A. Uniformization
Let be a non-compact Hermitian manifold. We generalize the notion of algebroid functions from to (see Definition 2.15). Let be a -valued algebroid function on Given a -leaf function element of When is not a singular branching point of we give a local Pusieux expansion of at i.e., in a local holomorphic coordinate we show that can be expanded locally into the following convergent Pusieux series (see Theorem 2.14)
at with where are meromorphic functions. By this Pusieux series, we construct a -leaf complex manifold and prove that
@@ B. Nevanlinna theory
The condition (2) is relaxed. For example, it holds when with and holds when with or and
@@
Assume that is non-positively curved satisfying
@@ C. Propagation of algebraic dependence
Assume that is either non-parabolic with non-negative Ricci curvature, or has non-positive sectional curvature satisfying (3).
@@
and (2) is satisfied if is non-parabolic with non-negative Ricci curvature, or (5) is satisfied if has non-positive sectional curvature satisfying (3).
@@ D. Unicity problems
Assume that is non-parabolic satisfying (2). Let be a positive integer. Consider the following condition
@@
Assume that is non-positively curved satisfying (3).
@@ 2.1. Notion of Algebroid Functions
Let be the indeterminacy locus of which is either an analytic set of complex codimension not smaller than for or an empty set for in Then, is holomorphic on
@@
The local definition for means that a collection of holomorphic functions i.e., is a holomorphic function for all satisfying that
@@ 2.1.3. Properties of function elements near a multiple point
According to Corollary 2.6, for any with there exist two groups of distinct meromorphic function elements at subordinate to equation (7), respectively. Apply Theorem 2.10, then for a curve with and we are able to extend analytically to along In other words, there is a permutation of such that can continue analytically to along respectively. The symmetry implies that can also continue analytically to along However, we should mention that won’t be, in general, continued analytically back to along respectively. In what follows, we would consider the analytic continuation of a function element along a closed Jordan curve surrounding any multiple point Take a smooth closed Jordan curve around such that which is close to and doesn’t surround or cross any connected components of that do not pass through Without loss of generality, we look at the analytic continuation of along starting from When returns to along it will turn to certain function If then one continues to extend analytically along in the same direction of rotation. Repeating this action, we shall see that cycles on and on periodically. Let be the smallest period such that can return to itself along all such curves Then, we get We call a branching point with order of for Repeatedly, it can be seen that is divided into finitely many groups. Let denote the number of groups. If the -th group has members, then
@@
where is the polydisc centered at 0 with polyradius in Setting where is the unit disc centered at 0 in the -plane. For any we have Set and write
We show that is a single-valued analytic function on Notice that is the branching order of we see that is holomorphic on because is so close to 0 that there exist no any branching points of in We only need to examine that is a removed singularity of To do so, it suffices to prove that is bounded near 0 due to Riemann’s analytic continuation. In fact, it follows from (7) and (9) that
@@
at Let be the set of all branching points of in Say that a branching point of is non-singular, if does not lie in a crossing of components of and singular otherwise. Assume that is not a singular branching point of Then, is smooth if is small enough. Taking a suitable if necessary. Since is bounded on and defines the branching order of at which is independent of the choice of one can deduce from and Riemann’s analytic continuation that there exists a biholomorphic mapping with
@@
According to the above arguments, we show that is a -valued analytic function defined in a neighborhood of We call a -leaf function element. If one cuts along a real hypersurface passing through then can separate into single-valued analytic components, and all these components can turn to each other when they continue analytically along a closed curve around in that is sufficiently close to By the local expansion of we deduce that is a -multiple, -valued branching point with order of
Assume that is not a singular branching point of Set and define
@@
To conclude, for each there are distinct multi-leaf function elements subordinate to equation (7), denoted by where is a -valued meromorphic function with and Hence, has branching order at with When we say that is a branching point with order of Moreover, if is not a singular branching point of then we deduce that can expand locally into the convergent Pusieux series at That is, we have shown that
@@
Let be a -leaf function element subordinate to equation If is not a singular branching point of then can be expanded locally into the following convergent Pusieux series
at in a local holomorphic coordinate such that and for all where is the branching set of and are meromorphic functions.
@@
According to the analytic continuation and uniqueness theorem, the above arguments show that equation (7) uniquely defines a -valued meromorphic function on outside the branching points of Let denote the set of all branching points of which is an analytic set in called the branching set of We generalize the notion of algebroid functions as follows.
@@
In a word, denote by the distinct single-valued meromorphic components of and given a -leaf function element subordinate to equation (7), with distinct single-valued components Then, there is a permutation of such that on
In the end, we provide an example to illustrate how to calculate the order of branching points of algebroid functions.
Example. Calculating the order of the -valued algebroid function
on at branching points
Since have single-valued components, respectively, we infer that has single-valued components; Since have single-valued components, respectively, we infer that has single-valued components. Thus, is a -valued algebroid function. Let stand for any closed Jordan curve in All the branching sets of are as follows
@@
Take such that is surrounded by Note that are the only two branching sets of that pass through Therefore, is not allowed to surround or pass through any other branching sets of except for in order to determine the order of For all such we for
Thus, the order of is determined by the order of branching point of and the order of branching point of It is
@@
Take such that is surrounded by Note that are the only two branching sets of that pass through Similarly, cannot surround or pass through any other branching sets of except for It implies that for
Thus, the order of is determined by the order of branching points of It is
@@
Take such that is surrounded by Since are the only three branching sets of that pass through we have for
Thus, the order of is determined by the order of branching points of respectively. It is
@@
Take such that is surrounded by Since are the only three branching sets of that passes through we have for
Thus, the order of is determined by the order of branching points of It is
@@ 2.2. Multi-leaf Complex Manifolds by Algebroid Functions
for some two single-valued components of respectively, where is the branching set of with Set
@@
A meromorphic function element is called a direct analytic continuation of a multi-leaf function element if both and in a neighborhood of for some single-valued component of where is the branching set of If so, one may regard that
@@
Next, we will equip with a complex structure such that is a complex manifold. Taking a -leaf function element Assume that is not a singular branching point of and is chosen suitably. Thanks to Theorem 2.14, there exists a biholomorphic mapping such that can be expanded into the Pusieux series
@@
Let be the set of all function elements satisfying that is a singular branching point of Moreover, denotes the set of all the corresponding singular branching points. It is clear that is an analytic set of complex codimension not smaller than Set
@@
Treat the singular part Let be the natural projection by mapping any function element to Note that is a generic -sheeted analytic covering such that
@@
is a non-compact -leaf complex manifold of complex dimension We call the complex manifold by
@@ 2.3. Zero and Pole Divisors of Algebroid Functions
Let be a -valued algebroid function on defined by equation Let be single-valued meromorphic components of As a matter of convenience, we write
@@
where denotes the number of -valued points of a function at Without loss of generality, suppose that are all single-valued components such that which take at and suppose that are all single-valued components such that which take at Then, we have By Vieta’s theorem, it yields that
Combining this with (11), we see that reaches the maximal order of a pole of Since we obtain
which means that ∎
@@
3. Uniformization of Algebroid Functions
This section works on the uniformization of algebroid functions on a non-compact Hermitian manifold. As shown in Section 2.2, we have constructed a multi-leaf complex manifold of a multi-valued algebroid function on a non-compact Hermitian manifold In fact, this realizes the uniformization of by liftting to a meromorphic function on via the natural projection such that In what follows, we shall provide another way towards the uniformization of algebroid functions without constructing a multi-leaf complex manifold. The main idea originates from W. Stoll [35].
@@ 3.1. Algebroid Reduction
Let be a non-compact complex manifold of complex dimension Let be a generic -sheeted analytic covering, by which one means a proper, surjective holomorphic mapping with branching set i.e., the set of all points such that is not locally biholomorphic at Note that is an empty set or an analytic set of pure complex codimension It is clear that
is a proper, surjective, local biholomorphic mapping of sheet number and thus a -sheeted analytic covering. By R. Remmert [27] (see Andreotti-Stoll [2] also), the exceptional set
@@
where for any Naturally, is embedded into by the identity mapping Thus, we obtain
@@ 3.2. Uniformization
In what follows, we consider the case that For a convenience, we write and in short. Take a basis of Then for any it has a unique representation
for some numbers It gives a natural isomorphism defined by
which yields that If then (13) defines an algebraic equation
@@
Let be a -valued algebroid function on Then, there exist a non-compact complex manifold of complex dimension a proper surjective holomorphic mapping of generic sheet number and a meromorphic function such that
@@
in the sense of isomorphisms. Further, the Vieta’s theorem leads to
@@
Let be -valued points of Then
@@
In conclusion, given a -leaf complex manifold by a -valued algebroid function on there exists the natural projection of generic sheet number given by which is a proper surjective holomorphic mapping. That is, we have shown that
Corollary 3.8.
Let be the -leaf complex manifold by a -valued algebroid function on Then, can lift to a meromorphic function via the natural projection such that
@@ 4.1. First Main Theorem
Next, we introduce Nevanlinna’s functions. Let be a meromorphic function. Denote by the Fubini-Study form on where is the homogeneous coordinate of with and
The characteristic function of for is defined by
Let The proximity function of for with respect to is defined by
@@
The counting function of for with respect to is defined by
Lemma 4.1 (Dynkin Formula, [1, 11, 12]).
@@
Applying Dynkin formula to we obtain
@@ 4.2. Second Main Theorem for Non-negative Ricci Curvature
Assume that is non-parabolic with non-negative Ricci Curvature. Fix a reference point and denote by the geodesic ball centered at with radius in Let denote the Riemannian volume of The non-parabolicity of means that
Thus, we have the unique minimal positive global Green function of for which can be written as
where is the heat kernel of on Let be the Riemannian distance function of from By Li-Yau’s estimate [18], there exist constants such that
@@ 4.2.1. Construction of and Nevanlinna’s functions
Since and then we conclude that is a precompact domain containing satisfying that and In fact, the family exhausts i.e., for any sequence with we have
@@
4.2.2. Gradient estimates of Green functions
@@
According to the definition of Green function for it is immediate that for
Since for then we obtain
@@ 4.2.3. Calculus Lemma and Logarithmic Derivative Lemma
Lemma 4.7 (Borel Lemma).
@@
Apply Borel’s lemma to the left hand side of the above equality twice: one is to and the other is to then we deduce that for any there exists a subset of finite Lebesgue measure such that
@@ 4.2.4. Second Main Theorem
holds for all outside where one used the fact that Combining (4.2.4) with (20), we have the theorem proved. ∎
@@ 4.3. Second Main Theorem for Non-positive Sectional Curvature
Assume that has non-positive sectional curvature satisfying (3). From the definition of curvatures, it is not difficult to see that In particular, we have if has constant sectional curvature Fix a reference point We may assume without loss of generality that is a pole of for a technical reason if necessary, since, otherwise, one can consider the universal covering As long as is equipped with the pull-back metric induced from via then has the same curvatures as ones of In fact, is a Cartan-Hadamard manifold. In the following, we denote by the Riemannian distance function of from on
@@ A. Setups for
Since as and we deduce that is a precompact domain containing in which exhausts Thus, the boundary of can be formulated as

### v11

@@ Abstract.
- 3 Uniformization of Algebroid Functions without Construction
@@ 1.1. Motivation
where are holomorphic functions without common zeros in An algebroid function is viewed as a special multi-valued analytic function, while which was first put forward, G. Darboux deemed that it is an important class of functions. To see that, for example, it is known that a -valued algebroid function defines a -leaf Riemann surface which is a -sheeted ramified analytic covering of A further fact states that can lift to a single-valued meromorphic function on via a natural projection such that which is called the uniformization of
@@ A. Uniformization
Let be a non-compact Hermitian manifold. We generalize the notion of algebroid functions from to (see Definition 2.15). Let be a -valued algebroid function on Given a -leaf function element of When is not a singular branch point of we give a local Pusieux expansion of at i.e., in a local holomorphic coordinate we prove that can be expanded locally into the following convergent Pusieux series (see Theorem 2.14)
at with where are meromorphic functions. By this Pusieux series, we construct a -leaf complex manifold and show that
@@ B. Nevanlinna theory
The condition (2) is relaxed. For example, it is satisfied when with or when with or and
@@
Assume that has curvature satisfying
@@ C. Propagation of algebraic dependence
Assume that is either non-parabolic with non-negative Ricci curvature, or of curvature satisfying (3).
@@
and that either (2) is satisfied if is non-parabolic with non-negative Ricci curvature, or (5) is satisfied if has curvature satisfying (3).
@@ D. Unicity problems
Assume that is non-parabolic with volume growth satisfying (2). Let be a positive integer. Consider the following condition
@@
Assume that has curvature satisfying (3).
@@ 2.1. Notion of Algebroid Functions
Let be the indeterminacy locus of which is either an analytic set of complex codimension not smaller than for or an empty set for in Note that is holomorphic on
@@
The local definition for means that i.e., a collection of holomorphic functions for all such that
@@ 2.1.3. Properties of function elements near a multiple point
According to Corollary 2.6, for any with there exist two groups of distinct meromorphic function elements at subordinate to equation (7), respectively. Apply Theorem 2.10, then for a curve with and we are able to extend analytically to along In other words, there is a permutation of such that can continue analytically to along respectively. The symmetry implies that can also continue analytically to along However, we should mention that won’t be, in general, continued analytically back to along respectively. In what follows, we would consider the analytic continuation of a function element along a closed Jordan curve surrounding any multiple point Take a smooth closed Jordan curve around such that which is close to and doesn’t surround or cross any connected components of that do not pass through Without loss of generality, we look at the analytic continuation of along starting from When returns to along it will turn to certain function by which we mean that for a local coordinate chart ( is small enough), jumps to some single-valued component when travels around once along started at due to the change of -dimensional argument of If then one continues to extend analytically along in the same direction of rotation. Repeat this action, we shall see finally that cycles on and on periodically. Let be the smallest period such that can return to itself along all such curves Then, we can obtain We call a branch point with order of if Repeatedly, it can be seen that is divided into finitely many groups. Let be the number of groups. If the -th group has members for then
@@
where is the polydisc centered at 0 with polyradius in Setting where is the unit disc centered at 0 in the -plane. For any we have
Set and write
We show that is a single-valued analytic function on Notice that is the branch order of we deduce that is holomorphic on because is so close to 0 that there exist no any branch points of in We only need to examine that is a removed singularity of To do so, it suffices to prove that is bounded near 0 due to Riemann’s analytic continuation. In fact, it follows from (7) and (9) that
@@
at Let be the set of all branch points of in Saying that a branch point of is non-singular, if does not lie in any crossings of components of and singular otherwise. Assume that is not a singular branch point of Then, is smooth when is small enough. Taking a suitable if necessary. Since is bounded on and defines the branch order of at that is independent of the choice of we can conclude from and Riemann’s analytic continuation that there exists a biholomorphic mapping with
@@
According to the above arguments, we show that is a -valued analytic function defined in a neighborhood of We call a -leaf function element. If one cuts along a real hypersurface passing through then can separate into single-valued analytic components, and all these components can turn to each other when they continue analytically along a closed curve around in that is sufficiently close to By the local expansion of we note that is a -multiple, -valued branch point with order of
Assume that is not a singular branch point of We set and define
@@
To conclude, for each there are distinct multi-leaf function elements subordinate to equation (7), denoted by where is a -valued meromorphic function with and Hence, has branch order at with If then we say that is a branch point with order of Furthermore, if is not a singular branch point of then we deduce that can expand locally into the convergent Pusieux series at That is, we have shown that
@@
Let be a -leaf function element subordinate to equation If is not a singular branch point of then can be expanded locally into the following convergent Pusieux series
at in a local holomorphic coordinate such that and for all where is the branch set of and are meromorphic functions.
@@
According to the analytic continuation and uniqueness theorem, the above arguments show that equation (7) uniquely defines a -valued meromorphic function on outside the branch points of Let stand for the set of all branch points of which is an analytic set in called the branch set of We extend the notion of algebroid functions from to as follows.
@@
In a word, denote by the distinct single-valued components of and given a -leaf function element subordinate to equation (7), with distinct single-valued components Then, there exists a permutation of with for on
In the end, we provide an example to illustrate how to calculate the order of branch points of algebroid functions.
Example. Calculating the branch order of the algebroid function
on at branch points
Since have single-valued components, respectively, we infer that has single-valued components; Since have single-valued components, respectively, we infer that has single-valued components. Thus, is a -valued algebroid function. Let stand for any closed Jordan curve in All the branch sets of are as follows
@@
Take such that is surrounded by Note that are the only two branch sets of that pass through So, is not allowed to surround or pass through any other branch sets of except for in order to determine the order of For all such we for
Hence, the order of is determined by the orders of branch point of and the order of branch point of It is
@@
Take such that is surrounded by Note that are the only two branch sets of that pass through Similarly, cannot surround or pass through any other branch sets of except for It implies that for
Hence, the order of is determined by the orders of branch points of It is
@@
Take such that is surrounded by Since are the only three branch sets of that pass through we have for
Hence, the order of is determined by the orders of branch points of respectively. It is
@@
Take such that is surrounded by Since are the only three branch sets of that passes through we have for
Hence, the order of is determined by the orders of branch points of It is
@@ 2.2. Multi-leaf Complex Manifolds by Algebroid Functions
for some two single-valued components of respectively, where is the branch set of with Set
@@
A meromorphic function element is called a direct analytic continuation of a multi-leaf function element if both and in a neighborhood of for some single-valued component of where is the branch set of If so, one may regard that
@@
Next, we will equip with a complex structure such that is a complex manifold. Taking a -leaf function element Assume that is not a singular branch point of Let be chosen suitably. By means of Theorem 2.14, there exists a biholomorphic mapping such that can be expanded into the Pusieux series
@@
Let stand for the set of all function elements such that is a singular branch point of Moreover, we denote by the set of all the corresponding singular branch points of which is obvious an analytic set of complex codimension not smaller than Set
@@
Finally, we shall treat the singular branch set Let be the natural projection which maps a function element to Note that is a -sheeted ramified analytic covering such that
@@
is a -sheeted ramified analytic covering of We call the -leaf complex manifold by
Corollary 2.21 (Uniformization).
An algebroid function on can lift to a meromorphic function on via the natural projection by mapping to such that
@@ 2.3. Zero and Pole Divisors of Algebroid Functions
Let be a -valued algebroid function on defined by equation Let be single-valued components of
@@
where denotes the number of -valued points of a function at Without loss of generality, we assume that are all components that take at According to Vieta’s theorem, we obtain
It implies that since reaches the maximal pole order at It yields from (11) that
Thus, we conclude that ∎
@@
3. Uniformization of Algebroid Functions without Construction
In this section, we provide another approach to realize the uniformization of algebroid functions without constructing finite ramified analytic coverings. Our main idea originates from W. Stoll [35].
@@ 3.1. Algebroid Reduction
Let be a non-compact complex manifold of complex dimension Let be a -sheeted ramified analytic covering, which means a proper, surjective holomorphic mapping with a ramification set i.e.,
is a proper, surjective, local biholomorphic mapping of sheet number and thus a -sheeted analytic covering. Here, is the set of those points such that is not locally biholomorphic at Note that is an empty set or an analytic set of pure complex codimension By R. Remmert [27] (see Andreotti-Stoll [2] also), the exceptional set
@@
where for any Naturally, is embedded into by the mapping Thus, we obtain
@@ 3.2. Uniformization
In what follows, we consider the case when For convenience, we write and for short. Take a basis of For any there is a unique representation
for certain It induces a natural isomorphism defined by
which follows that If then (13) defines an algebraic equation
@@
Let be an algebroid function. Then, there exist a non-compact complex manifold of complex dimension a proper surjective holomorphic mapping of generic sheet number and a meromorphic function such that
@@
in the sense of isomorphisms. In further, the Vieta’s theorem leads to
@@
Let be the -valued points of Then
@@ 4.1. First Main Theorem
Next, we introduce Nevanlinna’s functions. Let be a meromorphic function. Denote by the Fubini-Study form on where is the homogeneous coordinate of with and with
The characteristic function of on with respect to is defined by
Let The proximity function of on with respect to is defined by
@@
The counting function of on with respect to is defined by
We need the following Dynkin formula (see, e.g., [1, 11, 12]).
Lemma 4.1 (Dynkin Formula).
@@
Applying Dynkin formula to we get
@@ 4.2. Second Main Theorem for Non-negative Ricci Curvature
Assume that is non-parabolic. Fix a reference point and denote by the geodesic ball centered at with radius in Let denote the Riemannian volume of The non-parabolicity of means that
Thus, we have the unique minimal positive global Green function of for We can write as
in which is the heat kernel of on Let be the Riemannian distance function of from By Li-Yau’s estimate [18], there exist constants such that
@@ 4.2.1. Construction of and Nevanlinna’s functions
Since
we see that is a precompact domain containing with and Moreover, the family can exhaust i.e., for any sequence with we have
@@
4.2.2. Gradient estimate of Green functions
@@
Invoking the definition of Green function for it is immediate to deduce that for
Since for we obtain
@@ 4.2.3. Calculus Lemma and Logarithmic Derivative Lemma
Lemma 4.7 (Borel’s Growth Lemma).
@@
Applying Borel’s growth lemma to the left hand side of this equality twice: one is to and another is to we deduce that for any there exists a subset of finite Lebesgue measure such that
@@ 4.2.4. Second Main Theorem
holds for all outside in which we used the fact that Combining (4.2.4) with (20), we have the theorem proved. ∎
@@ 4.3. Second Main Theorem for Non-positive Sectional Curvature
Assume that has curvature satisfying (3). According to the definition of curvature, it is not difficult to see that In particular, we have when has constant sectional curvature Fix a reference point We may assume without loss of generality that is a pole of for a technical reason if necessary, since, otherwise, one can consider the universal covering If is equipped with pullback metric induced from via then preserves the same curvature as Note that is a Cartan-Hadamard manifold. In the following, we use to denote the Riemannian distance function of from on
@@ A. Setups for
Since as and one infers that is a precompact domain containing in which exhausts Thus, the boundary of can be formulated as